# The trimer paradox: the effect of stiff constraints on equilibrium distributions in overdamped dynamics

### v1

@@ Abstract
We reconsider the classical problem of a freely joined chain of Brownian particles connected by elastic springs and study its conformational probability distribution function in the overdamped regime in the limit of infinite stiffness of constraints. We show that the well-known solution by Fixman is missing a shape-related term, later alluded to but not computed by Helfand. In our approach, the shape term, also termed zero-point energy, arises explicitly from a careful treatment of the distributional limit. We present a computationally feasible method of calculation of the shape term and demonstrate its validity in a couple of examples.
@@ I Introduction
Molecular Dynamics and Brownian Dynamics simulations are now standard tools for detailed modelling of a plethora of molecular and mesoscopic systems, where structural complexity poses a challenge to theoretical calculations. Upon the introduction of a suitable coarse-graining scheme to represent the composition of a given molecule, its conformational space is prescribed by specifying intramolecular interactions, such as chemical bonds or electrostatic forces, between the subunits. The nature of these interactions endows molecular models with flexibility, which then in turn affects their conformational variability, as well as their macroscopic and statistical properties, such as rheology, diffusivity, and thermodynamics, and also biological function.
@@
A popular idealisation for a wide variety of systems is purely mechanical, where the molecule is represented by beads connected with springs, an idea which probably originated from Kramers [1], who built on concepts of Kuhn, and termed it the pearl-necklace model. Variations of this idea are now the cornerstone of polymer physics [2], but have also been used to describe other classical many-body systems. Physically, the introduction of intramolecular interactions leads to constrained dynamics with multiple time scales. Typically, the characteristic time scales associated with internal vibrations of molecular bonds are much shorter than those of translational and rotational motion or bond angle dynamics. The disparity of time scale of relaxation of hard (or stiff) degrees of freedom (such as bond lengths) in comparison to soft degrees of freedom (such as bond angles) is a source of stiffness in the problem and hinders fast numerical simulation of such systems [3].
@@
In many applications, the dynamics of the hard degrees of freedom are of secondary importance, and one method of circumventing this difficulty is to treat them as fully constrained (i.e. rigid), thus eliminating the short timescale, and enabling faster calculation. This is the basis of many algorithms, such as SHAKE [4] and its further extensions. During the development of such rigid models, a disconnect was discovered between the equilibrium bond angle distributions in the rigid simulations and the simulations with bonds modelled as very stiff springs.
@@
The simplest example in which the complexity of the constrained dynamics can be appreciated is the case of a flexible trimer: a hypothetical molecule containing 3 subunits and 2 harmonic bonds, depicted in Fig. 2a. The angle between the two bonds is denoted by $\psi$. In the classical textbook [3], Frenkel and Smit present an apparently paradoxical result: the marginal distribution of the bond angle in the case of rigid bonds $\mathrm{d}p_{\text{rigid}}$ and the limiting distribution of spring-like bonds where the spring stiffness $k$ is taken to infinity $\mathrm{d}p_{k\to\infty}$ do not coincide. More precisely, they report
$\mathrm{d}p_{k\to\infty}$ $=$ $N_{1}\sin\psi\,\,\mathrm{d}\psi$ (1)
$\mathrm{d}p_{\text{rigid}}$ $=$ $N_{2}\sin\psi\sqrt{{1-\frac{\cos^{2}\psi}{4}}}\,\,\mathrm{d}\psi$ (2)
@@
for appropriate normalisation constants $N_{1}$ and $N_{2}$. Qualitatively, distribution (1) corresponds to a spherically uniform distribution of the second bond direction when the coordinates are aligned with the first bond, while in Eq. (2) it seems that an additional force resists bond alignment. In an early and widely cited work, Fixman [5] presented a derivation to argue that the ratio of the two probability densities, Eqs. (1) and (2), can be computed with the knowledge of the constraining surfaces alone and does not require knowledge of the shape of the confining potential. This presupposes that both of these distributions are well defined whenever the constraint is given, which we argue is not the case.
@@
In this contribution, we present a clear mathematical procedure leading to the aforementioned limits in the overdamped regime when inertial effects can be neglected. We show that the distribution given by Eq. (1) is not completely determined by fixing the bond lengths. To correctly predict the limiting distribution, knowledge about the nature of confinement is needed and its structure survives in the final expression. Furthermore, we demonstrate that Fixman’s expressions for the stiff limit miss a term, which is important even in the case of harmonic springs.
@@
The first calculations of the equilibrium properties of flexible polymers can be traced to the papers by Kramers [1], Gō and Sheraga [6] and Fixman [5]. Results of the latter paper were used essentially without changes in works such as [7, 8, 9, 10]. Early papers focused on the treatment of constraints in the Lagrangian and Hamiltonian pictures, both classical and quantum-mechanical, by defining the soft and stiff degrees of freedom and integrating out generalised coordinates corresponding to the hard variables. The limiting distribution is then obtained for infinitely stiff constraints when stiff variables are eliminated.
@@
Fixman [5] presented a purely classical calculation of the limiting distribution, correctly writing the action of the distribution on a test function but erroneously assuming that the constraining potential depends only on the hard coordinates. As a result, details of the confining potential cannot appear in his final expressions, which contain only the projected phase-space volume element terms. Gō and Sheraga [6] treated the problem quantum-mechanically and correctly identified two contributions to the equilibrium probability density of the soft coordinates: the projected volume elements and the zero-point energy of the vibrational motion of the hard coordinates. However, they then stated that the zero-energy contribution in typical molecular systems varies only a little when changing the soft coordinates and may be altogether neglected. We show in the following that this is not necessarily the case in overdamped dynamics. By focusing on a concrete classical two-dimensional example with one constraint, Helfand [11] correctly identified the rigid-rod type distribution as a uniform distribution on the constraining manifold and computed the stiff-spring limit in a similar simple case. Unfortunately, these results are not expressed in a form applicable to higher-dimensional examples with multiple constraints.
@@
Due to this lack of generality, later works such as that of Hinch [8], cite Helfand as a reference to an erroneous claim that the bond angle distribution becomes non-uniform in the case of rigid-rod system. The pitfall lies in the assumption that the distribution should be uniform under an action of a rotation matrix on a single bond, presumably by analogy with a dimer. However, there is a distinct difference between these two cases; In the case of a dimer, the true physical symmetries of the equations of motion – rigid body rotation of the whole system and translations – act transitively on the space of all possible configurations, thus the equilibrium distribution can be determined from symmetry arguments. More generally, for a rigid body the rotation group $\mathrm{SO}(3)$ can be treated as a topological manifold, resulting in a natural notion of uniformity, with the uniform distribution given by the Haar measure. For the trimer, however, we have configurations that are meaningfully distinct. There is no symmetry that changes the bond angle $\psi$. While the equilibrium distribution for 6 out of the 7 degrees of freedom can be determined from the six symmetries of translation and rotation, the relative weight of configurations with different $\psi$ has to refer to the physical problem, not just geometric considerations.
@@
Hinch [8] approached the problem differently, by manipulating the Langevin equations, which are notoriously difficult to handle, and erroneously assumed that details of the confining potential are of little importance to the equilibrium distribution of the constrained configurations. Van Kampen and Lodder [12] noted that the approach of Helfand [11] is generally applicable to constrained systems and comment that the discrepancy between the stiff-spring and rigid-rod distributions in the case of the trimer molecule is due to the fact that the width of this gully is not the same everywhere [12]. However, they did not quantify the effect of this width in a higher-dimensional case. As a result, the knowledge of the influence of the shape of the constraining potential appears to have been lost as multiple works (for example, [13, 14, 3]) mention only the volume element densities and omit the zero point energy or, equivalently, the potential shape terms from their descriptions of constrained dynamics, widely using the result of Fixman [5] as reference.
@@ III The stiff spring limit
We present a derivation of the limiting distribution in the following setting, sketched in Fig. 1. Consider a molecule (or a polymer) with a handful of subunits (atoms, beads, or monomers) and $N$ degrees of freedom in a heat bath at constant temperature $T$, with $\beta=1/k_{B}T$. The conformation of the molecule is described by an $N$-dimensional vector $\bm{q}$. We parameterise conformations by Cartesian coordinates of monomers $q_{i}$ with $i\in 1\dots N$. The potential energy of the molecule has two components: the conformation-dependent energy $U(\bm{q})$ and the confining (springs) potential $k^{2}W(\bm{q})$ where $k$ is a large parameter describing the spring stiffness. Note that $k$ here would correspond to the square root of a harmonic spring constant. The probability density $p_{k}$ of the equilibrium distribution of the system configuration with respect to the Lebesgue measure $\mathrm{d}\bm{q}$ is given by the Boltzmann distribution
@@
with an appropriate normalising function $N(k)$. Suppose now that the spring potential is non-negative, $W(\bm{q})\geq 0$, and attains a minimum for a configuration lying on a smooth submanifold $\mathcal{M}$ of dimension $M<N$, so that $W(\bm{q})=0\iff\bm{q}\in\mathcal{M}$. To determine the weak limit of $p_{k}$, we consider integrals $I_{k}$ for a compactly supported smooth test function $\phi(\bm{q})$ given by
@@
Since $\mathcal{M}$ is smooth, we can use the tubular embedding theorem to define new orthonormal coordinates $\bm{q}\mapsto(\bm{r},\bm{s})=(r_{1},r_{2},\dots,r_{M},s_{1},s_{2},\dots)$ in the vicinity of $\mathcal{M}$, where $\bm{q}\in\mathcal{M}\iff\bm{s}=0$. In other words, coordinates $r_{i}$ represent the soft degrees of freedom and $s_{i}$ represent the hard directions. Since the new coordinates are orthonormal we know that the determinant of the transformation $(q_{i})\mapsto(r_{i},s_{j})$ equals 1. Additionally, $I_{k}$ is finite by compactness of the support of $\phi$ and continuity of $p_{k}$, so by Fubini’s theorem we can replace the volume integral by an iterated one. Then, $I_{k}$ is given by
$I_{k}=\iint p_{k}(\bm{r},\bm{s})\phi(\bm{r},\bm{s})\mathrm{d}\bm{s}\mathrm{d}\bm{r}$ (5)
@@
$I_{k}=\iint N(k)\exp\left\{-\beta U(\bm{r},\bm{s})-\beta k^{2}\left[\bm{s}^{T}\bm{\mathsf{H}}(\bm{r})\bm{s}+o(||\bm{s}||^{2})\right]\right\}\phi(\bm{r},\bm{s})\mathrm{d}\bm{s}\mathrm{d}\bm{r}$ (6)
@@
$I_{k}=\iint\frac{N(k)}{k}\exp\left\{-\beta U(\bm{r},k^{-1}\bm{t})-\beta\left[\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}+||\bm{t}||^{2}\frac{o(k^{-2}||\bm{t}||^{2})}{k^{-2}||\bm{t}||^{2}}\right]\right\}\phi(\bm{r},k^{-1}\bm{t})\mathrm{d}\bm{t}\mathrm{d}\bm{r}$ (7)
@@
$I=\iint\left(\lim_{k\to\infty}\frac{N(k)}{k}\right)\exp\left[-\beta U(\bm{r},\bm{0})-\beta\,\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}\right]\phi(\bm{r},\bm{0})\mathrm{d}\bm{t}\mathrm{d}\bm{r}$ (8)
@@
$I=L\int|\bm{\mathsf{H}}(\bm{r})|^{-1/2}\exp\left[-\beta U(\bm{r},\bm{0})\right]\phi(\bm{r},\bm{0})\mathrm{d}\bm{r}$ (9)
@@
$L=\lim_{k\to\infty}\frac{N(k)\sqrt{(2\pi)^{N-M}}}{k},$ (10)
@@
III.1 Computation of the shape term
A careful choice of parametrisation allows for computation of the metric term with relative ease. However, the Hessian term $|\bm{\mathsf{H}}|$ may be more difficult to compute. In this Section, we present a feasible approach to its computation.

### v2

@@ Abstract
We reconsider the classical problem of a freely joined chain of Brownian particles connected by elastic springs and study its conformational probability distribution function in the overdamped regime in the limit of infinite stiffness of constraints. We show that the well-known solution by Fixman [Proc. Natl. Acad. Sci. USA 71, 3050 (1974)] is missing a shape-related term, later alluded to but not computed by Helfand [J. Chem. Phys 71, 5000 (1979)]. In our approach, the shape term, also termed zero-point energy, arises explicitly from a careful treatment of the distributional limit. We present a computationally feasible method for the calculation of the shape term and demonstrate its validity in a couple of examples.
@@ I Introduction
Molecular Dynamics (MD) and Brownian Dynamics (BD) simulations are now standard tools for detailed modelling of a plethora of molecular and mesoscopic systems, where structural complexity poses a challenge to theoretical calculations. Upon the introduction of a suitable coarse-graining scheme to represent the composition of a given molecule, its conformational space is prescribed by specifying intramolecular interactions, such as chemical bonds or electrostatic forces, between the subunits. The nature of these interactions endows molecular models with elasticity, which in turn affects their conformational variability, as well as their macroscopic and statistical properties, such as rheology, diffusivity, and thermodynamics, and also biological function.
@@
A popular idealisation for a wide variety of systems is purely mechanical, where the molecule is represented by beads connected with springs, an idea which probably originated from Kramers [1], who built on concepts of Kuhn, and termed it the pearl-necklace model. Variations of this idea are now the cornerstone of polymer physics [2], but have also been used to describe other classical many-body systems. Physically, the introduction of intramolecular interactions leads to constrained dynamics with multiple time scales. Typically, the characteristic time scales associated with internal vibrations of molecular bonds are much shorter than those of translational and rotational motion or bond angle dynamics. The disparity in the time scale of relaxation of hard (or stiff) degrees of freedom (such as bond lengths) compared to soft degrees of freedom (such as bond angles) is a source of stiffness in the problem and hinders fast numerical simulation of such systems [3].
@@
In many applications, the dynamics of the hard degrees of freedom are of secondary importance, and one method of circumventing this difficulty is to treat them as fully constrained (i.e. rigid), thus eliminating the short timescale, and enabling faster simulations. This is the basis of many algorithms, such as SHAKE [4] and its further extensions. During the development of such rigid models, a disconnect was discovered between the equilibrium bond angle distributions in the rigid simulations and the simulations with bonds modelled as very stiff springs.
@@
The simplest example in which the complexity of the constrained dynamics can be appreciated is the case of a flexible trimer: a hypothetical molecule containing 3 subunits and 2 harmonic bonds, depicted in Fig. 2a. The angle between the two bonds is denoted by $\psi$. In the classical textbook [3], Frenkel and Smit present an apparently paradoxical result: the marginal distribution of the bond angle in the case of rigid bonds, $\mathrm{d}p_{\text{rigid}}$, and the limiting distribution of spring-like bonds where the spring stiffness $k$ is taken to infinity, $\mathrm{d}p_{k\to\infty}$, do not coincide. More precisely, they report
$\mathrm{d}p_{k\to\infty}$ $=$ $N_{1}\sin\psi\,\,\mathrm{d}\psi,$ (1)
$\mathrm{d}p_{\text{rigid}}$ $=$ $N_{2}\sin\psi\sqrt{{1-\frac{\cos^{2}\psi}{4}}}\,\,\mathrm{d}\psi,$ (2)
@@
for appropriate normalisation constants $N_{1}$ and $N_{2}$. Qualitatively, the distribution (1) corresponds to a spherically uniform distribution of the second bond direction when the coordinates are aligned with the first bond, while in Eq. (2) it seems that an additional force resists bond alignment. In an early and widely cited work, Fixman [5] presented a derivation to argue that the ratio of the two probability densities, Eqs. (1) and (2), can be computed with the knowledge of the constraining surfaces alone and does not require knowledge of the shape of the confining potential. This presupposes that both of these distributions are well defined whenever the constraint is given, which we argue is not the case.
@@
In this contribution, we present a clear mathematical procedure leading to the aforementioned limits in the overdamped regime when inertial effects can be neglected. We show that the distribution given by Eq. (1) is not completely determined by fixing the bond lengths. To correctly predict the limiting distribution, knowledge about the nature of confinement is needed, and its structure survives in the final expression. Furthermore, we demonstrate that Fixman’s expressions for the stiff limit miss a term, which is important even in the case of harmonic springs.
@@
The first calculations of equilibrium properties of flexible polymers can be traced to the papers of Kramers [1], Gō and Sheraga [6] and Fixman [5]. The results of the latter paper were used essentially without changes in later works, such as [7, 8, 9, 10]. Early papers focused on the treatment of constraints in the Lagrangian and Hamiltonian pictures, both classical and quantum-mechanical, by defining the soft and stiff degrees of freedom and integrating out generalised coordinates corresponding to the hard (stiff) variables. The limiting distribution is then obtained for infinitely stiff constraints and the stiff variables are eliminated.
@@
Fixman [5] presented a purely classical calculation of the limiting distribution, correctly writing the action of the distribution on a test function but erroneously assuming that the constraining potential depends only on the hard coordinates. As a result, the details of the confining potential cannot appear in his final expressions, which contain only the projected phase-space volume element terms. Gō and Sheraga [6] treated the problem quantum-mechanically and correctly identified two contributions to the equilibrium probability density of the soft coordinates: the projected volume elements and the zero-point energy of the vibrational motion of the hard coordinates. However, they then stated that the zero-energy contribution in typical molecular systems varies only a little when changing the soft coordinates and may be altogether neglected. We show in the following that this is not necessarily the case in overdamped dynamics. By focusing on a concrete classical two-dimensional example with one constraint, Helfand [11] correctly identified the rigid-rod type distribution as a uniform distribution on the constraining manifold and computed the stiff-spring limit in a similar simple case. Unfortunately, these results are not expressed in a form applicable to higher-dimensional examples with multiple constraints.
@@
Due to this lack of generality, later works cite Helfand as a reference to an erroneous claim that the bond angle distribution becomes non-uniform in the case of a rigid-rod system. The pitfall lies in the assumption that the distribution should be uniform under the action of a rotation matrix on a single bond, presumably by analogy with a dimer. However, there is a distinct difference between these two cases; In the case of a dimer, the true physical symmetries of the equations of motion – rigid body rotation of the whole system and translations – act transitively on the space of all possible configurations, and thus the equilibrium distribution can be determined from symmetry arguments. More generally, for a rigid body, the rotation group $\mathrm{SO}(3)$ can be treated as a topological manifold, resulting in a natural notion of uniformity, with the uniform distribution given by the Haar measure. For the trimer, however, we have configurations that are meaningfully distinct. There is no symmetry that changes the bond angle $\psi$. Although the equilibrium distribution for 6 out of 7 degrees of freedom can be determined from the six symmetries of translation and rotation, the relative weight of configurations with different $\psi$ has to refer to the physical problem, not just geometric considerations.
@@
Hinch et al [8, 12] approached the problem differently, by manipulating the Langevin equations, which are notoriously difficult to handle, and erroneously assumed that details of the confining potential are of little importance to the equilibrium distribution of the constrained configurations. Van Kampen and Lodder [13] noted that the approach of Helfand [11] is generally applicable to constrained systems and comment that the discrepancy between the stiff-spring and rigid-rod distributions in the case of the trimer molecule is due to the fact that the width of this gully is not the same everywhere [13]. However, they did not quantify the effect of this width in a higher-dimensional case. As a result, the knowledge of the influence of the shape of the constraining potential appears to have been lost as multiple works (for example, [14, 15, 3]) mention only the densities of the volume elements and omit the zero point energy or, equivalently, the potential shape terms from their descriptions of constrained dynamics, widely using the result of Fixman [5] as a reference.
@@ III The stiff spring limit
This work concerns two distributions – the rigid rod distribution and the distributional limit of equilibrium distributions at increasingly stiff confinement: the stiff spring limit. We begin our consideration with a careful definition of the stiff spring limit with the rigid rod distribution discussed in later sections.
We present a derivation of the limiting distribution in the following setting, sketched in Fig. 1. Consider a molecule (or a polymer) with a handful of identical subunits (atoms, beads, or monomers) each of mass $m$ and a total of $2N$ degrees of freedom ($N$ configurational and $N$ momentum) in a heat bath at constant temperature $T$, with $\beta=1/k_{B}T$. The conformation of the molecule is described by an $N$-dimensional vector $\bm{q}$ and the velocities of monomers are specified by a vector $\bm{v}$. We parametrise the conformations by Cartesian coordinates of monomers $q_{i}$ and Cartesian components of velocities $v_{i}$ with $i\in 1\dots N$. The potential energy of the molecule has two components: the conformation-dependent energy $U(\bm{q})$ and the confining (springs) potential $k^{2}W(\bm{q})$ where $k$ is a large parameter describing the spring stiffness. Note that $k$ here would correspond to the square root of a harmonic spring constant. The probability density $\hat{p}_{k}$ of the equilibrium distribution of the system configuration with respect to the Lebesgue measure $\mathrm{d}\bm{q}\mathrm{d}\bm{v}$ is given by the Maxwell-Boltzmann distribution [16]
$p_{k}=\hat{N}(k)\exp\left\{-\beta\left[U(\bm{q})+k^{2}W(\bm{q})+\frac{1}{2}m\bm{v}\cdot\bm{v}\right]\right\},$ (3)
with an appropriate normalising function $\hat{N}(k)$. In the case under consideration $m$ does not depend on $\bm{q}$ so the momentum coordinates are easy to integrate out and we arrive at the familiar Boltzman probability density $p_{k}$ with respect to the measure $\mathrm{d}\bm{q}$ [16] written as
@@
with a (different) normalising function $N(k)$. Crucially, this density is given with respect to canonical conformation coordinates and not just any coordinates – if generalised coordinates are used, more care is needed and volume element density terms appear in expressions analogous to Eq. (4). Moreover, in the case of generalised coordinates, the expression for kinetic energy is also more complicated. Clearly, Eq. (4) is a physical object and we are free to work in any coordinates of choice, and for the purpose of the derivation of the stiff spring limit Cartesian coordinates are particularly convenient. We return to the treatment of generalized coordinates in later chapters of this manuscript.
With these caveats outlined, we define the stiff spring limit $\mathrm{d}p_{k\to\infty}$ as a measure which satisfies
$\mathrm{d}p_{k\to\infty}=\lim_{k\to\infty}\mathrm{d}p_{k},$ (5)
in the weak sense of the limit.
Suppose now that the spring potential is non-negative, $W(\bm{q})\geq 0$, and attains a minimum for a configuration lying on a smooth submanifold $\mathcal{M}$ of dimension $M<N$, so that $W(\bm{q})=0\iff\bm{q}\in\mathcal{M}$. To determine the weak limit of $p_{k}$, we consider integrals $I_{k}$ for a compactly supported smooth test function $\phi(\bm{q})$ given by
@@
Since $\mathcal{M}$ is smooth, we can use the tubular embedding theorem to define new orthonormal coordinates $\bm{q}\mapsto(\bm{r},\bm{s})=(r_{1},r_{2},\dots,r_{M},s_{1},s_{2},\dots)$ in the vicinity of $\mathcal{M}$, where $\bm{q}\in\mathcal{M}\iff\bm{s}=0$. In other words, coordinates $r_{i}$ represent the soft degrees of freedom and $s_{i}$ represent the hard directions. Since the new coordinates are orthonormal, we know that the determinant of the transformation $(q_{i})\mapsto(r_{i},s_{j})$ equals 1. Additionally, $I_{k}$ is finite by the compactness of the support of $\phi$ and the continuity of $p_{k}$, so by Fubini’s theorem we can replace the volume integral by an iterated one. Then, $I_{k}$ is given by
$I_{k}=\iint p_{k}(\bm{r},\bm{s})\phi(\bm{r},\bm{s})\mathrm{d}\bm{s}\mathrm{d}\bm{r}.$ (7)
@@
$I_{k}=\iint N(k)\exp\left\{-\beta U(\bm{r},\bm{s})-\beta k^{2}\left[\bm{s}^{T}\bm{\mathsf{H}}(\bm{r})\bm{s}+o(||\bm{s}||^{2})\right]\right\}\phi(\bm{r},\bm{s})\mathrm{d}\bm{s}\mathrm{d}\bm{r}.$ (8)
@@
$I_{k}=\iint\frac{N(k)}{k}\exp\left\{-\beta U(\bm{r},k^{-1}\bm{t})-\beta\left[\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}+||\bm{t}||^{2}\frac{o(k^{-2}||\bm{t}||^{2})}{k^{-2}||\bm{t}||^{2}}\right]\right\}\phi(\bm{r},k^{-1}\bm{t})\mathrm{d}\bm{t}\mathrm{d}\bm{r}.$ (9)
@@
$I=\iint\left(\lim_{k\to\infty}\frac{N(k)}{k}\right)\exp\left[-\beta U(\bm{r},\bm{0})-\beta\,\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}\right]\phi(\bm{r},\bm{0})\mathrm{d}\bm{t}\mathrm{d}\bm{r},$ (10)
@@
$I=L\int|\bm{\mathsf{H}}(\bm{r})|^{-1/2}\exp\left[-\beta U(\bm{r},\bm{0})\right]\phi(\bm{r},\bm{0})\mathrm{d}\bm{r},$ (11)
@@
$L=\lim_{k\to\infty}\frac{N(k)\sqrt{(2\pi\beta)^{N-M}}}{k},$ (12)
@@
III.1 Calculation of the shape term
Figure 2: (a) Parametrisation of the trimer uses the position of the central particle ($b$), bond angle $\psi$ and bond lengths $l_{a}$ and $l_{b}$, with identical equilibrium spring lengths $l_{0}$. Rigid-body rotation around the particle $b$ is omitted for clarity. (b) Motion along the symmetric (top) and antisymmetric eigenvectors of the hessian $\bm{\mathsf{H}}$ of the harmonic confinement function. Note that the central bead must move for the overall motion to be orthogonal to the constraining manifold $\mathcal{M}$. (c) Equilibrium probability density function (PDF) for the bond angle $\psi$ from the Langevin numerical simulation. Solid line is the theoretical prediction of Eq. (34).
A careful choice of parametrisation allows one to calculate the metric term with relative ease. However, the calculation of the Hessian term $|\bm{\mathsf{H}}|$ may be more difficult. In this Section, we present a feasible approach to its determination.